OpenStudy (anonymous):

Should there be a decimal trig functions? ex...... cotA = .29

OpenStudy (campbell_st):

why not cot = 1/tan so cot(a) = 0.29 which means tan(A) = 100/29

OpenStudy (campbell_st):

but the question doesn't make a lot of sense unless you are talking about exact values

OpenStudy (jdoe0001):

\(\bf cot(A) = 0.29\implies \cfrac{1}{tan(A)} = 0.29\implies \cfrac{1}{0.29}= tan(A)\)

OpenStudy (jdoe0001):

and that gives you a valid angle

OpenStudy (jdoe0001):

what's the context?